1. What are the vector components of the following: Round off your answer to two decimal numbers.

a. G = 10m, 1500°

b. L = 10m, 900°

c. A = 10m, 3600°

d. D = 10m, 360° N of W


1
Expert's answer
2021-09-28T11:14:32-0400

"a.\space G_x= 10\sin(1500°)=\\\space\space\space\space\space\space\space\space\space\space=10\sin(360°·4+60)=10\sin60°=8.7\text{ m},\\
\space \space\space\space G_y=10\cos(1500°)=\\\space\space\space\space\space\space\space\space\space\space=10\cos(360°·4+60)=10\cos60°=5.0\text{ m}.\\\space\\
b.\space L_x= 10\sin(900°)=\\\space\space\space\space\space\space\space\space\space\space=10\sin(360°·2+180°)=0\text{ m},\\
\space \space\space\space L_y=10\cos(900°)=\\\space\space\space\space\space\space\space\space\space\space=10\cos(360°·2+180°)=-10\text{ m}.\\\space\\
c.\space A_x= 10\sin(3600°)=\\\space\space\space\space\space\space\space\space\space\space=10\sin(360°·10+0)=0\text{ m},\\
\space \space\space\space A_y=10\cos(3600°)=\\\space\space\space\space\space\space\space\space\space\space=10\cos(360°·10+0)=10\text{ m}.\\\space\\

d.\space D_x= 10\sin(-360°+180°)=0\text{ m},\\
\space \space\space\space D_y=10\cos(-360°+180°)=-10\text{ m}.\\\space\\"